Use the factoring to solve the quadratic equation. Check by substitution or by using a graphing utility.
5x^2 =8x+21

Answer by mananth(16946)

5x^2 =8x+21
5x^2 -8x-21=0
5x^2-15x+7x-21=0
5x(x-3)+7(x-3)=0
(x-3)(5x+7)=0
x=3 OR x=-7/5
5*3^2=45
8*3+21=45